# On the Sequence Evaluation based on Stochastic Processes

### v3

@@ Abstract
Generative models have gained significant prominence in Natural Language Processing (NLP), especially in tackling the complex task of modeling and evaluating long text sequences. This task is crucial for advancing various downstream applications, such as text generation and machine translation. Recent methods that utilize stochastic processes to capture the intrinsic dynamics of sequences have shown superior performance in generative modeling. However, the accurate encoding of both temporal and structural dependencies from text datasets, as well as leveraging this encoded information for sequence evaluation, remains an open area of research. In this paper, we propose a novel approach to learn the stochastic dynamics of long text sequences, utilizing a negative log-likelihood-based encoder that outperforms contrastive learning methods. We also introduce a likelihood-based evaluation metric for long-text assessment, which measures sequence coherence and can be applied to downstream tasks such as Human-AI discrimination. Our encoder preserves sequence coherence effectively and performs robustly on out-of-domain datasets. Additionally, the proposed evaluation metric captures both temporal and structural information comprehensively. Theoretical analysis demonstrates the superiority of our metric in sequence evaluation, and experimental results highlight its flexibility and exceptional performance across a variety of tasks, showcasing its utility in diverse NLP applications.
@@ 1 Introduction
Generative models are rapidly gaining traction in NLP (Zou et al., 2023; Yang et al., 2023; Yi et al., 2024), particularly in tackling the intricate task of modeling and generating long text sequences. This challenge is critical for various downstream applications, including text generation and machine translation. Recently, the use of stochastic representations to model latent spaces has emerged as a promising solution, demonstrating considerable potential in diverse fields such as time-series analysis (Liu et al., 2021), dynamical flow modeling (Albergo et al., 2023; Albergo & Vanden-Eijnden, 2023), and video generation (Zhang et al., 2023). In the domain of text generative models, Wang et al. (2022) proposed a method that frames long sequences as stochastic dynamical flows, achieving competitive results in generating coherent long texts. However, the challenge of accurately learning the time-dependent probability density functions embedded within text datasets remains unresolved. Moreover, fully harnessing the information encoded in the stochastic representation continues to pose a significant challenge that has yet to be comprehensively addressed.
Brownian Bridge for stochastic representation modeling and coherence metric. Linguistic theory offers valuable insights into modeling long articles as stochastic processes. For example, coherent long texts tend to exhibit certain common properties: (1) a central theme that remains consistent throughout the text, (2) sentences at the beginning and end that emphasize this central idea, and (3) sentences in the middle that may diverge slightly from the theme but still remain aligned with it. (Sheng et al., 2024) These characteristics naturally align with the concept of a “bridge” in stochastic processes, where the structure supports a consistent trajectory while allowing for controlled variability. In this work, we employ the Brownian Bridge (abbreviated as BB) to represent this conceptual bridge. It is a special stochastic process characterized by “fixed” starting and ending positions (Øksendal & Øksendal, 2003) and has been widely applied in various fields. Inspired by the work of Sheng et al. (2024) and Wang et al. (2022), we discover that the fit of a trajectory as a BB can be used to evaluate and represent several characteristics of long-form text as an evaluation metric and text encoder.
@@
Figure 1: Schematic diagram of the Stochastic Representation and Metric. An article from domain ${\mathcal{X}}_{i}$, segmented into sentences $(u_{1},u_{2},\cdots u_{n})$, is processed by the Stochastic Process (or Contrastive Learning) encoder, which consists of a pre-trained large language model (LLM) and a multi-layer perceptron (MLP). The LLM generates a raw representation $(x_{1},x_{2},\cdots x_{n})$, which the MLP transforms into a time-dependent latent trajectory $(s_{1},s_{2},\cdots s_{n})$. Additionally, we introduce a stochastic process metric (SPM) that captures the temporal and structural properties of this latent trajectory.
@@
First, we introduce a novel likelihood-based evaluation metric, Stochastic Process Metric (SPM) for long-text assessment that quantifies coherence by leveraging both temporal and structural dependencies of the stochastic process, with detailed explanation in Section 2.1. Unlike similar approaches that rely heavily on the training domain (Lai & Tetreault, 2018; Jeon & Strube, 2022) and are limited by their training schemes – which typically evaluate only pairwise data and restrict comparisons to articles of the same length, SPM offers greater flexibility while maintaining comparable performance. To demonstrate this, we generalize the established Shuffle test (Barzilay & Lapata, 2005; Joty et al., 2018; Moon et al., 2019) into a Mixed Shuffle test. Instead of merely comparing shuffled and unshuffled versions of the same article, this new test also compares across the shuffle and unshuffled versions between different articles, which can be used to evaluate the robustness of the metric regardless of individual article properties, e.g. article length. Moreover, SPM proves useful in downstream tasks such as Human-AI discrimination and performs robustly in out-of-domain (O.O.D.) tasks, likely due to it capture the structural information of the stochastic representation.
Second, we present a novel text encoder, Stochastic Process Encoder (SP Encoder) for effectively capturing the stochastic dynamics present in extensive collections of long articles. We propose a negative log-likelihood-based training objective for text encoders, comparing its effectiveness to that of contrastive learning methods (Wang et al., 2022). The SP Encoder translates sequences into a generalized BB while retaining essential information, leading to improved performance on O.O.D. tests. In contrast, while contrastive learning can be useful for assessing coherence, it may fall short in adequately capturing the representation of the target stochastic process, which could hinder its effectiveness on other tasks, particularly those involving O.O.D. data.
In summary, the main contributions of our work are as follows:
  • We propose a novel likelihood-based metric (SPM), designed to capture both temporal and structural dependencies with a solid theoretical foundation. The SPM demonstrates superior flexibility and effectiveness in sequence evaluation, as evidenced by its performance across various tasks.
@@
  • We hypothesized and validated that fitting the stochastic representation, as measured by the SPM, can serve as an effective and flexible metric for multiple downstream tasks, such as distinguishing between human and AI-generated writing.
  • We introduced a negative log-likelihood-based method (SP Encoder), for learning stochastic representations that preserves sequence coherence while efficiently utilizing latent dimensions. Compared to contrastive learning encoders, it exhibits robust and superior performance across O.O.D. datasets.
@@ 2.1 Stochastic representation
In this section, we introduce a stochastic representation of the encoded sequences by modeling them using BBs, which are commonly utilized stochastic processes in various applications. We begin by defining a standard BB $\{B(t):t\in[0,T]\}$ with $B(0)=0$ and $B(T)=0$. For any $t\in[0,T]$, the process $B(t)$ follows a normal distribution $B(t)\sim N(0,t(T-t)/T)$. Additionally, for $s,t\in[0,T]$ with $s<t$, the covariance between $B(s)$ and $B(t)$ is given by ${\rm Cov}(B(s),B(t))=s(T-t)/T$. A general BB can then be constructed as $a+(t/T)(b-a)+\sigma B(t)$, where $a$ and $b$ are fixed start and end points, respectively, and $\sigma$ is the standard deviation of the process.
@@
For an input sequence $\bar{{\mathbf{s}}}=(s_{0},\ldots,s_{T})$ with $s_{t}\in\mathbb{R}^{d}$ for $t=0,1,\ldots,T$, we capture temporal dependence using standard BBs. To account for structural dependence among the components, we consider $d$ independent standard BBs $B_{1}(t),\ldots,B_{d}(t)$ over the interval $[0,T]$. At each time $t$, the sequence is modeled as $s_{t}=\mu_{t}+{\mathbf{W}}(B_{1}(t),\ldots,B_{d}(t))^{\top}$, where ${\mathbf{W}}\in{\mathbb{R}}^{d\times d}$ is a transformation matrix and $\mu_{t}=s_{0}+(t/T)(s_{T}-s_{0})$ represents the mean at time $t$. The structural dependence is captured by $\Sigma={\mathbf{W}}{\mathbf{W}}^{\top}$. Let ${\mathbf{s}}=(s_{1},\ldots,s_{T-1})$ denote the sequence excluding the start and end points, and let $\bm{\mu}=(\mu_{1},\ldots,\mu_{T-1})$ be the corresponding means.
@@
The SP Encoder and the SPM are based on the likelihood function of the input sequences, with $\Sigma$ being the only unknown parameter. The following proposition presents the likelihood function.
@@
Let $\Sigma_{T}\in{\mathbb{R}}^{(T-1)\times(T-1)}$ be the covariance matrix with entries $[\Sigma_{T}]_{s,t}=s(T-t)/T$. For $n$ independent input sequences $\bar{\mathbf{s}}_{1},\ldots,\bar{\mathbf{s}}_{n}$ with lengths $T_{1}+1,\ldots,T_{n}+1$, generated by the same ${\mathbf{W}}$ (or equivalently, $\Sigma$), the log-likelihood function is
$\ell(\Sigma|\{\bar{{\mathbf{s}}}_{i}\}_{i=1}^{n})=$ $\frac{d\sum_{i=1}^{n}(T_{i}-1)}{2}\log(2\pi)-\frac{d}{2}\sum_{i=1}^{n}\log(\lvert\Sigma_{T_{i}}\rvert)-\frac{\sum_{i=1}^{n}(T_{i}-1)}{2}\log(\lvert\Sigma\rvert)$
$-\frac{1}{2}\sum_{i=1}^{n}\mathrm{tr}(\Sigma^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})\Sigma_{T_{i}}^{-1}({\mathbf{s}}_{i}-\bm{\mu}_{i})^{\top}).$
By Proposition 1, we derive the log-likelihood function given the input sequences, which enables us to compute the maximum likelihood estimate (MLE) of $\Sigma$.

### v4

# BBScoreV2: Learning Time-Evolution and Latent Alignment from Stochastic Representation

@@ Abstract
Autoregressive generative models play a key role in various language tasks, especially for modeling and evaluating long text sequences. While recent methods leverage stochastic representations to better capture sequence dynamics, encoding both temporal and structural dependencies and utilizing such information for evaluation remains challenging. In this work, we observe that fitting transformer-based model embeddings into a stochastic process yields ordered latent representations from originally unordered model outputs. Building on this insight and prior work, we theoretically introduce a novel likelihood-based evaluation metric BBScoreV2. Empirically, we demonstrate that the stochastic latent space induces a "clustered-to-temporal ordered" mapping of language model representations in high-dimensional space, offering both intuitive and quantitative support for the effectiveness of BBScoreV2. Furthermore, this structure aligns with intrinsic properties of natural language and enhances performance on tasks such as temporal consistency evaluation (e.g., Shuffle tasks) and AI-generated content detection.
BBScoreV2: Learning Time-Evolution and Latent Alignment from Stochastic Representation
Tianhao Zhang1thanks: Equal contribution., Zhecheng Sheng [^1] footnotemark: 1, Zhexiao Lin211footnotemark: 1, Chen Jiang [^1] footnotemark: 1, Dongyeop Kang1 1University of Minnesota, Twin Cities 2University of California, Berkeley {zhan7594, sheng136, jian0649, dongyeop}@umn.edu, zhexiaolin@berkeley.edu
@@ 1 Introduction
Generative models are rapidly gaining traction in NLP (Zou et al., 2023; Yang et al., 2023; Yi et al., 2024), particularly for the complex task of modeling and generating long text sequences—a challenge central to downstream applications such as text generation and machine translation. Recently, stochastic representations of latent spaces have emerged as a promising approach, showing considerable success in areas including time-series analysis (Liu et al., 2021), dynamical flow modeling (Albergo et al., 2023; Albergo and Vanden-Eijnden, 2023), and video generation (Zhang et al., 2023). In the context of text generation, Wang et al. (2022) introduced a method that models long sequences as stochastic dynamical flows, yielding strong results in producing coherent long texts. However, accurately learning the time-dependent probability density functions inherent in text data remains an open problem. Furthermore, effectively leveraging the information encoded in stochastic representations continues to be a significant challenge that has not yet been fully addressed.
Brownian bridge (BB) process helps to learn time-evolution in the stochastic representation
While the temporal evolution captured in articles offers insights into linguistic properties like coherence and theme (Sheng et al., 2024), effectively encoding this temporal information into latent representations remains difficult. Drawing inspiration from the Time-control model (Wang et al., 2022) and Stochastic Interpolation (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023), we propose using the "bridge process" from stochastic process theory (Øksendal and Øksendal, 2003) to encode and evaluate sentence-level temporal information within latent representations. Furthermore, by leveraging the raw embeddings from frozen language models, we can also incorporate sentence-level structural information. In this work, we utilize the BB, the simplest bridge process characterized by fixed start and end points (Øksendal and Øksendal, 2003) and widely applied across various domains. We believe that more complex bridge processes, such as the Schrödinger bridge (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023), could offer richer encoding capabilities, representing a promising avenue for future research.
@@
Figure 1: Schematic diagram of the Stochastic Representation in the latent space. An article from domain ${\mathcal{X}}_{i}$, segmented into sentences $(u_{1},u_{2},\cdots u_{n})$, is processed by the encoder which consists of a pre-trained language model (LM) and a multi-layer perceptron (MLP). The encoder maps each sentence into latent space and after optimizing for the stochastic objective, the latent trajectory becomes time dependent.
@@
To evaluate such encoded time-evolution information, we introduce BBScoreV2, a novel likelihood-based evaluation metric for long-text assessment. BBScoreV2 evaluates the time evolution within a stochastic representation by considering both its temporal and structural dependencies, as detailed in Section 3.1. This metric is particularly useful for article coherence evaluation, exemplified by the shuffle task, which disrupts the natural temporal order. Existing methods for this task (Lai and Tetreault, 2018; Jeon and Strube, 2022) often depend heavily on the training domain, are limited by their training paradigms, can only assess pairwise data, and are restricted to articles of the same length. In contrast, BBScoreV2 assesses general temporal order, offering greater flexibility while maintaining comparable performance. To demonstrate this, we generalize the standard Shuffle test (Barzilay and Lapata, 2005; Joty et al., 2018; Moon et al., 2019) into a more robust Mixed Shuffle test. This new test compares shuffled and unshuffled versions both within and between different articles, allowing for evaluation of the metric’s robustness independent of individual article characteristics like length. Furthermore, BBScoreV2 proves valuable in downstream applications such as Human-AI discrimination and exhibits strong performance in out-of-domain (O.O.D.) scenarios, likely due to its ability to capture the general structural and temporal information in human writing and preserve it in the stochastic representation.
The main contributions of our work can be summarized as follows:
  • We demonstrate that clustered language model embeddings can be effectively structured into temporal ordered stochastic representations via a simple multi-layer architecture.
  • We propose a novel likelihood-based metric (BBScoreV2) to evaluate temporal and structural dependencies within the stochastic representation with solid theoretical foundation.
@@
  • We hypothesized and validated that temporal and structural information encoded in the stochastic representation, as measured by the BBScoreV2, can potentially serve as an effective and flexible metric for multiple downstream tasks such as coherence evaluate and AI-generated text detection.
2 Related work
Stochastic processes have demonstrated robust capabilities in modeling complex tasks across various fields, including biology (Horne et al., 2007) and finance (Øksendal and Øksendal, 2003). Recently, the use of stochastic representations to model latent spaces has shown considerable promise in diverse applications such as time-series analysis (Liu et al., 2021) and dynamical flow modeling (Albergo et al., 2023; Albergo and Vanden-Eijnden, 2023). Notably, such methods also excel in generation tasks, including video generation (Zhang et al., 2023), and long text generation (Wang et al., 2022). A critical aspect of these tasks is to incorporate time-evolution into the latent representation, which requires capturing the time-dependent probability density functions embedded within real-world data. Generally, there are two approaches to tackle this challenge. One method is the likelihood-free training paradigm (Durkan et al., 2020), exemplified by contrastive learning techniques, which have demonstrated significant effectiveness in handling high-dimensional data (van den Oord et al., 2018; Wang et al., 2022; Zhang et al., 2023). This approach enables the learning of predictive density indirectly, rather than through direct reconstruction (Mathieu et al., 2021). The alternative method is the traditional likelihood-based approach, such as stochastic interpolants (Albergo et al., 2023; Albergo and Vanden-Eijnden, 2023), which requires the pre-definition of specific target stochastic processes. Both methods exhibit substantial potential in their respective tasks.
Coherence of articles, as defined by Reinhart (1980), referring to the logical flow and connection of ideas in a text, is one of the most complex temporal dynamic encoded in the articles. Studies have shown that transformers, while effective in generating tasks, often struggle with capturing coherence (Deng et al., 2022). To improve how language models learn long-text dynamics, methods using latent spaces have been developed (Bowman et al., 2016; Gao et al., 2021), focusing on sentence embeddings by considering neighboring utterances. However, these methods often produce static representations and neglect the text’s dynamic nature. A recent approach using stochastic representations, such as the BB, incorporates "temporal dynamics" to improve long-range text dependencies (Wang et al., 2022). This method shows promise in generating coherent long texts through capturing structural and temporal information.
In addition to generative tasks, evaluating coherence in a given text also remains a challenge (Sheng et al., 2024; Maimon and Tsarfaty, 2023). Building on stochastic concepts, Sheng et al. (2024) developed a heuristic metric for coherence assessment, grounded in the unsupervised learning approach proposed by Wang et al. (2022). This score demonstrated considerable performance on artificial shuffle tasks. However, their method relies on a heuristic understanding of the BB and fails to adequately establish a theoretical foundation for the metric setup, which limit the effectiveness and flexibility of their score, particularly its sensitivity to article length.
@@ 3.1 Brownian bridge process
In this section, we introduce a stochastic representation of the encoded sequences by modeling them using BBs. We begin by defining a standard BB $\{B(t):t\in[0,T]\}$ with $B(0)=0$ and $B(T)=0$. For any $t\in[0,T]$, the process $B(t)$ follows a normal distribution $B(t)\sim N(0,t(T-t)/T)$. Additionally, for $s,t\in[0,T]$ with $s<t$, the covariance between $B(s)$ and $B(t)$ is given by ${\rm Cov}(B(s),B(t))=s(T-t)/T$. A more general BB start from $a$ and end at $b$ can then be constructed as $a+(t/T)(b-a)+\sigma B(t)$, where $a$ and $b$ are fixed start and end points, respectively, and $\sigma$ is the standard deviation of the process.
3.2 Contrastive learning encoder
The encoder architecture consists of two components: a frozen, pre-trained language model and a trainable multilayer perceptron (MLP) network. We extract the hidden state corresponding to the end-of-sentence (EOS) token from the last layer of the language model. This hidden state serves as an input to a four-layer MLP, which is trained to map the input to the latent space. The purpose of the encoder is to learn a non-linear mapping from the raw input space to the latent space, denoted as $f_{\theta}:{\mathcal{X}}\to{\mathcal{S}}$. We train the encoder using contrastive learning (CL) loss ($L_{\rm CL}$), which enhances its ability to differentiate between positive and negative samples, following the approach of (van den Oord et al., 2018; Wang et al., 2022).
We adopt the CL encoder framework as presented by Wang et al. (2022). In this framework, a key structural assumption is imposed on the latent space, namely an isotropic covariance structure represented by $\Sigma={\mathbf{I}}_{d}$, where ${\mathbf{I}}_{d}$ denotes the $d$-dimensional identity matrix. Consequently, for an arbitrary starting point ${\mathbf{s}}_{0}$ at time $t=0$ and an ending point ${\mathbf{s}}_{T}$ at time $t=T$, the marginal distribution of ${\mathbf{s}}_{t}$ at time $t$ is given by Equation 1.
[…]
Consider any triplet of observations $({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3})$ with ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}\in{\mathcal{X}}$. The goal is to ensure that $f_{\theta}({\mathbf{x}}_{2})$ follows the above marginal distribution with starting point $f_{\theta}({\mathbf{x}}_{1})$ and ending point $f_{\theta}({\mathbf{x}}_{3})$. For a sequence of observations $({\mathbf{x}}_{0},\ldots,{\mathbf{x}}_{T})$, let $B=\{({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T})\}$ be a batch consisting of randomly sampled positive triplets $({\mathbf{x}}_{0},{\mathbf{x}}_{t},{\mathbf{x}}_{T})$ with $0<t<T$. Then, the CL loss function $L_{\rm CL}$ is defined by Equation 2.
To further investigate the structural assumption ($\Sigma={\mathbf{I}}_{d}$) employed during encoder training, particularly given the importance of latent space correlation structure for downstream tasks, we conducted ablation studies. Specifically, we tested two different encoders: 1) CL encoder with AnInfoNCE loss: a CL loss designed by Rusak et al. (2024) to keep learning the covariance matrix $\Sigma$ during training, and 2) a negative-log-likelihood based method (SP Encoder) which is purely based on fitting the temporal distribution of the bridge process.
3.3 Alignment in latent space
@@
To evaluate trajectories within the stochastic latent space, we propose a method to approximate the inherent correlation structure and assess both spatial and temporal properties of the encoded latents. For an input sequence $\bar{{\mathbf{s}}}=(s_{0},\ldots,s_{T})$ with $s_{t}\in\mathbb{R}^{d}$ for $t=0,1,\ldots,T$, we capture temporal dependence using standard BBs. To account for structural dependence among the components, we consider $d$ independent standard BBs $B_{1}(t),\ldots,B_{d}(t)$ over the interval $[0,T]$. At each time $t$, the sequence is modeled as $s_{t}=\mu_{t}+{\mathbf{W}}(B_{1}(t),\ldots,B_{d}(t))^{\top}$, where ${\mathbf{W}}\in{\mathbb{R}}^{d\times d}$ is a transformation matrix and $\mu_{t}=s_{0}+(t/T)(s_{T}-s_{0})$ represents the mean at time $t$. The structural dependence is captured by $\Sigma={\mathbf{W}}{\mathbf{W}}^{\top}$. Let ${\mathbf{s}}=(s_{1},\ldots,s_{T-1})$ denote the sequence excluding the start and end points, and let $\boldsymbol{\mu}=(\mu_{1},\ldots,\mu_{T-1})$ be the corresponding means.
@@
The proposed BBScoreV2 is based on the likelihood function of the input sequences, with $\Sigma$ being the only unknown parameter. The following proposition presents the likelihood function. For the detailed proof, please check Appendix C
@@
Let $\Sigma_{T}\in{\mathbb{R}}^{(T-1)\times(T-1)}$ be the covariance matrix with entries $\Sigma_{T}]_{s,t}=s(T-t)/T$. For $n$ independent input sequences $\bar{\mathbf{s}}_{1},\ldots,\bar{\mathbf{s}}_{n}$ with lengths $T_{1}+1,\ldots,T_{n}+1$, generated by the same $\mathbf{W}$} (or equivalently, $\Sigma$), and then the log-likelihood function is defined in Equation 3.
By Proposition 1, given the input sequences, the maximum likelihood estimate (MLE) of $\Sigma$ is.